OpenStudy (anonymous):

b^2 + 2b + 1 factored ?

OpenStudy (imstuck):

You are wanting to factor this?

OpenStudy (anonymous):

i got that one i need a new one now, can you help me factor a^2 - a -6

OpenStudy (anonymous):

(b+1)(b+1) or (b+1)^2, it is also called a perfect square I believe.

OpenStudy (anonymous):

Yes !! thats what i got :)

OpenStudy (anonymous):

a^2 - a - 6 (a-3)(a+2)

OpenStudy (anonymous):

Do you want some help so you know how to figure them out on your own?

OpenStudy (anonymous):

I can teach you basic factoring methods...

OpenStudy (anonymous):

yes please could you show me on this one n^2 + 8n - 9

OpenStudy (anonymous):

Well how I would show you basic factoring is a little different than starting with this specific example, but it seems you are needing help with homework so, n^2 + 8n - 9 -> (n+9)(n-1)

OpenStudy (anonymous):

Always look at your last component which is C in your standard form (Ax^2+Bx+C) of a linear equation. Ask yourself, what two #'s do I multiply to get this number? In the case of 9 previous, we can do 9*1, 3*3, and that's it.

OpenStudy (anonymous):

Standard form actually includes Ax^2+Bx+C=0.

OpenStudy (anonymous):

okay well i need those to find out common denominators. i know how to do the reallllly simple ones but i just dont think hard enough for those so i need the answers quick and easy. the other one was n^2 - 1 so i just need to combine like terms

OpenStudy (anonymous):

so (n-1) (n+1) ( n+9) (n-1)

OpenStudy (anonymous):

and combine it

OpenStudy (anonymous):

n^2-1 is a difference of squares, (n-1)(n+1)

OpenStudy (anonymous):

and what would it look like if i combined it ?

OpenStudy (anonymous):

Sorry, not a difference of squares... not sure. If you combined (n-1)(n+1) and ( n+9) (n-1)?

OpenStudy (anonymous):

yes!!

OpenStudy (anonymous):

would it be like (n-1) (n+1) (n+9) ?

OpenStudy (anonymous):

You would come to a 4th degree polynomial. What is the problem asking exactly?

OpenStudy (anonymous):

" find at least common denominator for these two rational expressions"

OpenStudy (anonymous):

okay well the answer is (n-1) (n+1) (n+9) !!

OpenStudy (anonymous):

okay new one a^2 - 2a - 15

OpenStudy (anonymous):

(a-5)(a+3)

OpenStudy (anonymous):

Ok so a better answer is, Ask yourself which two numbers multiply to 15 that when you take the difference will = -2, which is your B component. You ask yourself the "difference" question when you have two (-)'s in your polynomial.

OpenStudy (anonymous):

3 and 5 multiply to 15, and 5-3 = 2

OpenStudy (anonymous):

Since B is negative, or B=-2, you make the larger number (5), the negative part (a-5)

OpenStudy (anonymous):

The A, B, C numbers, B and C in particular, are the Ax^2 + (B)x + (C) in your form. a^2 - 2a - 15 where A=1, B =-2, and C = -15, are you seeing what I'm saying?

OpenStudy (anonymous):

3x^2+4x+15 has A=3, B=4, and C=15 This is not factor-able so don't worry. when you have x^2, or any number ^2 without a leading coefficient, x^2, a^2, c^2... Your A=1. So x^2 +2x+1, A=1, B=2, and C=1

OpenStudy (anonymous):

Leading coefficient refers to the number before the x, a, or whatever is the first part. x^2, 2x^2, 3x^2: Your leading coefficient is 1, 2, and 3; or A= 1, 2, and 3

OpenStudy (anonymous):

So in your case your dealing with just the integer 1 as your leading coefficient for these polynomials you are needing to facvor.

OpenStudy (anonymous):

A=1, so we are not concerned with A at all. If your equation is +, +, or B and C are both positive: Ex. x^2 +2x+1, then your factors become (x+something)(x+something)

OpenStudy (anonymous):

Am I helping you at all?